Opened 12 years ago

Closed 11 years ago

#83 closed flight processing (fixed)

GB2007/06, flight day 192b/2007, New Forest

Reported by: mggr Owned by: mggr
Priority: alpha 4 high Milestone: 2007 data processing completion
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by mggr)

Data location: ~arsf/arsf_data/in_progress/2007/flight_data/uk/GB2007_06-2007_192b_New_Forest

Data arrived from ARSF via second SATA disk transfer at the start of Sept.

Scientific details: See ~arsf/arsf_data/in_progress/2007/ARSF_Applications-GB_2007/GB07-06_*

PI: Doreen Boyd

Sensors:

  • CASI (delivered 7/Dec/2007)

Initially this was thought to be a lidar only flight due to a permissions problem with transferring the CASI datasets. Created as a job in Nov.

Change History (16)

comment:1 Changed 12 years ago by anee

All lines processed fine, minor timing issues to fix.

comment:2 Changed 12 years ago by anee

Line 6 GPTOFF ~0.03, Line 7 GPTOFF ~0.03

comment:3 Changed 12 years ago by mggr

Vectors available at /users/rsg/arsf/vectors/from_akw/new_forest

comment:4 Changed 12 years ago by anee

Processed flight data located in: /users/rsg/anee/scratch_space/GB2007_06-2007_192b_New_Forest

Delivery directory created and filled.

comment:5 Changed 12 years ago by anee

  • Owner changed from anee to mggr

comment:6 Changed 12 years ago by anee

  • Owner changed from mggr to benj

comment:7 Changed 12 years ago by benj

Checked delivery files, minor changes to readme. Ready for delivery

comment:8 Changed 12 years ago by benj

  • Owner changed from benj to anee

comment:9 Changed 12 years ago by anee

CASI data delivered 07/Dec/2007.

comment:10 Changed 12 years ago by anee

  • Owner changed from anee to mggr

Needs to be archived.

comment:11 Changed 12 years ago by mggr

  • Description modified (diff)
  • Status changed from new to assigned

Synced into main repo at ~arsf/arsf_data/in_progress/2007/flight_data/uk/GB2007_06-2007_192b_New_Forest/

PI (Doreen) contacted us as she expected to have 15 bands but only had 14. This appears to be a configuration issue in flight (need to follow up).

comment:12 Changed 11 years ago by mggr

  • Resolution set to fixed
  • Status changed from assigned to closed

comment:13 Changed 11 years ago by mggr

Prepared for archiving (DTG, NEODC) - moved to ~arsf/arsf_data/complete-ready_to_archive/

comment:14 Changed 11 years ago by mggr

  • Resolution fixed deleted
  • Status changed from closed to reopened

(reopening until archived)

comment:15 Changed 11 years ago by mggr

Data sent to NEODC for archiving.

comment:16 Changed 11 years ago by mggr

  • Resolution set to fixed
  • Status changed from reopened to closed
Note: See TracTickets for help on using tickets.